Required information

[The following information applies to the questions displayed below.]

In 2021, the Westgate Construction Company entered into a contract to construct a road for Santa Clara County for $10,000,000. The road was completed in 2023. Information related to the contract is as follows:

2021 2022 2023
Cost incurred during the year $ 2,604,000 $ 4,032,000 $ 1,940,400
Estimated costs to complete as of year-end 5,796,000 1,764,000 0
Billings during the year 2,040,000 4,596,000 3,364,000
Cash collections during the year 1,820,000 4,000,000 4,180,000


Westgate recognizes revenue over time according to percentage of completion.

4. Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years assuming the following costs incurred and costs to complete information. (Do not round intermediate calculations and round your final answers to the nearest whole dollar amount. Loss amounts should be indicated with a minus sign.)

2021 2022 2023
Costs incurred during the year $ 2,604,000 $ 3,820,000 $ 3,220,000
Estimated costs to complete as of year-end 5,796,000 3,120,000 0

2021 2022 2023 Revenue Gross profit (loss)
Answer #1
Revenue recognition as per Percentage of Completion Method - Westgate Cosntruction Company
Year Actual cost incurred (A) Total cost incurred till date (B) Total estimated Cost ( C ) % of completion (D= B/C) Contract Price
(E )
Total Revenue to be recoganized
(F=E * D)
Revenue for current period (G) Gross Profit (H=G-A)
2021 $2,604,000 $2,604,000 $8,400,000 31.00% $10,000,000 $3,100,000 $3,100,000 $496,000
(2604000+5796000)
2022 $4,032,000 $6,636,000 $8,400,000 79.00% $10,000,000 $7,900,000 $4,800,000 $768,000
(6636000+1764000) (7900000-3100000)
2023 $1,940,400 $8,576,400 $8,576,400 100.00% $10,000,000 $10,000,000 2100000 $159,600
(10000000-7900000)